# Value-oriented forecast reconciliation for renewables in electricity markets

### v1

@@ 2.2 Aggregated Wind Energy Trading
Consider a scenario where there are $m$ WPPs, managed by a PM responsible for trading in the market on behalf of all the WPPs. Let $Y_{t+k,\texttt{sum}}$ denote the aggregate sum of the random variables $Y_{t+k,i}$, expressed as $Y_{t+k,\texttt{sum}}=\sum_{i}^{m}Y_{t+k,i}$. The observed value is indicated by $y_{t+k,\texttt{sum}}=\sum_{i}^{m}y_{t+k,i}$. Meanwhile, $y^{\texttt{F}}_{t+k,\texttt{sum}}$ stands for the trading offer by the PM, and $c_{t+k}(y_{t+k,\texttt{sum}}^{\texttt{F}},y_{t+k,\texttt{sum}})$ denotes the total imbalance cost associated with the offer $y^{\texttt{F}}_{t+k,\texttt{sum}}$. When the PM aims to minimize the total expected cost imbalance, the problem still reduces to a newsvendor, and the optimal offer of the portfolio is given by
$y_{t+k,\texttt{sum}}^{\texttt{F}*}=F^{-1}_{y_{t+k,\texttt{sum}}}(\alpha_{t+k}),$ (9)
where $F_{y_{t+k,\texttt{sum}}}$ denotes the CDF of $Y_{t+k,\texttt{sum}}$. The imbalance cost for the offer $y_{t+k,\texttt{sum}}^{\texttt{F}*}$ is then given as $c_{t+k}(y_{t+k,\texttt{sum}}^{\texttt{F}*},y_{t+k,\texttt{sum}})$. Accordingly, the expected total imbalance cost is calculated as $\mathbb{E}_{Y_{t+k,\texttt{sum}}}[c_{t+k}(y_{t+k,\texttt{sum}}^{\texttt{F}*},y_{t+k,\texttt{sum}})]$. The cost-saving benefits of aggregated trading, as demonstrated by Eppen, (1979) and Hartman et al., (2000), are outlined below.
@@
[Eppen, (1979); Hartman et al., (2000)] In a multi-location newsvendor system, the total optimal expected cost of the centralized system is lower than that of the decentralized system.
@@
$\mathbb{E}_{Y_{t+k,\texttt{sum}}}[c_{t+k}(y_{t+k,\texttt{sum}}^{\texttt{F}*},y_{t+k,\texttt{sum}})]\leq\sum_{i}^{m}\mathbb{E}_{Y_{t+k,i}}[c_{t+k}(y_{t+k,i}^{\texttt{F}*},y_{t+k,i})],$ (10)
which implies that offering as an aggregation can always lead to a reduction in costs for participants. Equality in (10) holds only when the variables $Y_{t+k,1},\cdots,Y_{t+k,m}$ are perfectly positively correlated. The aggregated wind energy offering problem has been explored by Bitar, (2011); Baeyens et al., (2013), which demonstrated that WPPs can consistently increase their expected profits through aggregated trading. However, these studies do not specify the assigned offer or the cost for each WPP. In most cases, the total of all independent offers does not align with the optimal offer of the PM. Using simple cost allocation rules, such as dividing the cost evenly among all WPPs, may leave some WPPs unsatisfied with their assigned costs.
@@
To address this, Freire et al., (2014); Nguyen and Le, (2018) have developed algorithms to identify allocation rules in the core of collaborative wind energy offering games. Then, the PM offers in the markets and allocates profits (or equivalently, costs) to WPPs based on the established allocation rule. The operational framework for the PM is shown in Figure 1. However, these studies often assume that both the PM and WPPs have access to identical information and that WPPs unanimously agree on the offers. In reality, diversified perspectives on future energy generation may lead to different opinions among the PM and the WPPs regarding the offers. Specifically, each WPP typically uses a forecast model to estimate its own independent offer, as shown in Equation (4). Similarly, the PM employs a forecasting model $f$ with parameters $\phi$ to predict the aggregated offer based on the feature vector $\mathbf{x}_{t}$. This model is expressed as
$y^{\texttt{F}}_{t+k,\texttt{sum}}=\hat{y}_{t+k,\texttt{sum}|t}=f(\mathbf{x}_{t};\phi).$ (11)
As the forecasts $\left\{\hat{y}_{t+k,1|t},\ldots,\hat{y}_{t+k,m|t},\hat{y}_{t+k,\texttt{sum}|t}\right\}$ are generated independently by the PM and WPPs, they are likely to be incoherent. Thus, WPPs may disagree with the offer proposed by the PM, and furthermore, the forecasts cannot be directly used to determine each WPP’s share in the submitted offer. We address this challenge by developing a value-oriented forecast reconciliation approach in Section 3.
@@ 3.1 Problem Formulation
For ease of reference, the following concise notations are introduced: $\left[y_{t+k,1},\ldots,y_{t+k,m},y_{t+k,\texttt{sum}}\right]^{\top}$ is represented by $\mathbf{y}_{t+k}$, while $\left[\hat{y}_{t+k,1|t},\cdots,\hat{y}_{t+k,m|t},\hat{y}_{t+k,\texttt{sum}|t}\right]^{\top}$ is expressed as $\hat{\mathbf{y}}_{t+k|t}$, and referred to as the base forecast. As shown in Figure 1, the interaction between WPPs and the PM inherently establishes a two-level hierarchical structure. In the context of forecast reconciliation (Athanasopoulos et al.,, 2024), forecasts at the leaf nodes of the hierarchy are referred to as bottom-level forecasts, while forecasts at other nodes are called aggregated forecasts. In our specific WPP-PM hierarchy, we can decompose the vector $\mathbf{y}_{t}$ into $m$ bottom-level values $\mathbf{b}_{t}\in\mathbb{R}^{m}$ and an aggregated value $y_{t,\texttt{sum}}$ such that $\mathbf{y}_{t}=\left[\mathbf{b}_{t}^{\top},y_{t,\texttt{sum}}\right]^{\top}$. By defining an aggregation matrix $\mathbf{S}\in\{0,1\}^{(m+1)\times m}$, we obtain
$\mathbf{y}_{t}=\mathbf{S}\mathbf{b}_{t}\Leftrightarrow\begin{bmatrix}\mathbf{b}_{t}\\ y_{t,\texttt{sum}}\end{bmatrix}=\begin{bmatrix}\mathbf{I}_{m}\\ \mathbf{s}_{\texttt{sum}}\end{bmatrix}\mathbf{b}_{t},\ \forall t,$ (12)
where $\mathbf{s}_{\texttt{sum}}=[1,1,\cdots,1]$ is a summation vector and $\mathbf{I}_{m}$ is the $m\times m$ identity matrix. Equivalently, formula (12) can be rewritten as
@@
where $\mathbf{B}=[-\mathbf{s}_{\texttt{sum}},1]$. This indicates that $\mathbf{y}_{t}$ lies within the null space of $\mathbf{B}$, expressed as $\mathbf{y}_{t}\in\texttt{null}(\mathbf{B})$. With the hierarchical structure, the coherent forecast is formally defined as follows.
@@
A forecast $\hat{\mathbf{y}}_{t+k|t}$ is said to be coherent iff $\hat{\mathbf{y}}_{t+k|t}\in\texttt{null}(\mathbf{B})$.
As previously stated, the base forecast, $\hat{\mathbf{y}}_{t+k|t}$ is independently generated by WPPs and PM, and is likely not coherent. Therefore, it is essential to reconcile the forecasts before using them in market trading. Let $\tilde{\mathbf{y}}_{t+k|t}=\left[\tilde{y}_{t+k,1|t},\ldots,\tilde{y}_{t+k,m|t},\tilde{y}_{t+k,\texttt{sum}|t}\right]^{\top}$ denote the reconciled forecasts, where
$\tilde{y}_{t+k,\texttt{sum}|t}=\sum_{i=1}^{m}\tilde{y}_{t+k,i|t}.$ (14)
@@
$\tilde{\mathbf{y}}_{t+k|t}=f_{\texttt{RE}}(\hat{\mathbf{y}}_{t+k|t}),$ (15)
where $f_{\texttt{RE}}$ represents the reconciliation function. In other words, we transform incoherent base forecast $\hat{\mathbf{y}}_{t+k|t}$ into coherent forecast $\tilde{\mathbf{y}}_{t+k|t}$ via the reconciliation function.
After reconciliation, the PM then places offers in the market using the reconciled aggregate forecast, $\tilde{y}_{t+k,\texttt{sum}|t}$. In parallel, the reconciled forecast for each WPP, $\tilde{y}_{t+k,i|t}$, is not submitted to the market; it serves instead to determine cost (or profit) allocation and is thus termed a pseudo-offer. Different from Section 2 where each WPP shares the same cost function, here the WPPs face heterogeneous cost functions, since they are apportioned as shares of the overall balancing cost. Concretely, the allocated cost for WPP $i$ is calculated based on the reconciled forecasts $\tilde{\mathbf{y}}_{t+k|t}$ and realizations $\mathbf{y}_{t+k}$, which is represented as $c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$. The explanation of $c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ is postponed to Section 3.2. Correspondingly, the profit for WPP $i$ in the context of aggregated wind energy trading, is represented as $\rho_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$, which can be expressed as
$\rho_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})=\pi^{\texttt{F}}_{t+k}y_{t+k,i}-c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k}).$ (16)
It is observed that, after reconciliation, the balancing cost of the aggregated forecast is always lower than the sum of costs of independent pseudo-offers, as described below.
Theorem 3.1.
Let $\left\{\tilde{y}_{t+k,1|t},\ldots,\tilde{y}_{t+k,m|t}\right\}$ be a set of $m$ independent pseudo-offers for the trading hour $t+k$. Then, almost surely,
$c_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}})\leq\sum_{i=1}^{m}c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i}),$
where $\tilde{y}_{t+k,\texttt{sum}|t}=\sum_{i=1}^{m}\tilde{y}_{t+k,i|t}$.
@@
Notably, achieving coherence is often effectively managed through existing reconciliation methods. In fact, simple approaches, such as bottom-up (BU) reconciliation, can successfully attain coherence. However, most existing quality-oriented forecast reconciliation methods assume that all agents will agree on the reconciliation procedure (Athanasopoulos et al.,, 2024). As discussed in Section 1, the challenge lies in achieving consensus on the reconciliation approach when agents have heterogeneous cost functions. Prior to formulating our reconciliation method, we present the following assumptions, reserving the specific methodologies for Section 3.2.
@@
In this context, we exclude sub-coalitions, as WPPs will only engage with the PM. Each WPP will decide whether to enter into an agency contract with the PM. Even if some WPPs choose to opt out of the contract, the PM can still manage the portfolio consisting of the remaining WPPs.
@@
This assumption is valid because WPPs are primarily focused on the individual benefits they derive from the coalition, rather than on the collective benefits of the coalition.

### v2

@@ 2.2 Aggregated Wind Energy Trading
Leveraging the complementary nature of geographically diverse renewable energy sources to hedge against price and quantity uncertainty in electricity markets has attracted significant interest (bitar2011selling). Consider a scenario where there are $m$ WPPs, managed by a PM responsible for trading in the market on behalf of all the WPPs. Let $Y_{t+k,\mathrm{sum}}$ denote the aggregate sum of the random variables $Y_{t+k,i}$, expressed as $Y_{t+k,\mathrm{sum}}=\sum_{i}^{m}Y_{t+k,i}$. The observed value is indicated by $y_{t+k,\mathrm{sum}}=\sum_{i}^{m}y_{t+k,i}$. Meanwhile, $y^{\mathrm{F}}_{t+k,\mathrm{sum}}$ stands for the trading offer by the PM, and $c_{t+k}(y_{t+k,\mathrm{sum}}^{\mathrm{F}},y_{t+k,\mathrm{sum}})$ denotes the total imbalance cost associated with the offer $y^{\mathrm{F}}_{t+k,\mathrm{sum}}$. When the generation assets are owned by the same company, the portfolio manager (PM) aims to minimize the total expected imbalance cost. In this case, the problem reduces to a classical newsvendor formulation, and the optimal offer for the aggregated portfolio is given by
$y_{t+k,\mathrm{sum}}^{\mathrm{F}*}=F^{-1}_{y_{t+k,\mathrm{sum}}}(\alpha_{t+k}),$ (9)
where $F_{y_{t+k,\mathrm{sum}}}$ denotes the CDF of the aggregated random variable $Y_{t+k,\mathrm{sum}}$. The imbalance cost associated with the offer $y_{t+k,\mathrm{sum}}^{\mathrm{F}*}$ is then given by $c_{t+k}(y_{t+k,\mathrm{sum}}^{\mathrm{F}*},y_{t+k,\mathrm{sum}})$. Accordingly, the expected total imbalance cost is computed as $\mathbb{E}_{Y_{t+k,\mathrm{sum}}}[c_{t+k}(y_{t+k,\mathrm{sum}}^{\mathrm{F}*},y_{t+k,\mathrm{sum}})]$.
The cost-saving benefits of aggregated trading—originally demonstrated by eppen1979note and further supported by hartman2000cores—are outlined below.
@@
[eppen1979note; hartman2000cores] In a multi-location newsvendor system, the total optimal expected cost of the centralized system is lower than that of the decentralized system.
@@
$\mathbb{E}_{Y_{t+k,\mathrm{sum}}}[c_{t+k}(y_{t+k,\mathrm{sum}}^{\mathrm{F}*},y_{t+k,\mathrm{sum}})]\leq\sum_{i}^{m}\mathbb{E}_{Y_{t+k,i}}[c_{t+k}(y_{t+k,i}^{\mathrm{F}*},y_{t+k,i})],$ (10)
which implies that offering as an aggregation can always lead to a reduction in costs for participants. Equality in (10) holds only when the variables $Y_{t+k,1},\cdots,Y_{t+k,m}$ are perfectly positively correlated.
In fact, bitar2011selling and baeyens2013coalitional have demonstrated that WPPs can consistently increase their expected profits through aggregated trading. As discussed above, in the centralized context, it is optimal to submit $y_{t+k,\mathrm{sum}}^{\mathrm{F}*}$ as the trading offer, as the primary objective is to maximize total profits. In contrast, in a multi-agent setting where the generation assets are owned by different companies, each agent seeks to maximize its own profit. As a result, the aggregate of independently determined offers often diverges from the PM’s optimal offer. This discrepancy necessitates determining both the total trading offer submitted to the market and the corresponding profit allocation for each individual agent. Applying simple cost allocation rules—such as evenly splitting the total cost among WPPs—may lead to dissatisfaction among participants due to perceived unfairness. To address this issue, various cooperative game-theoretic approaches have been employed to allocate costs among participating agents, including proportional sharing, the Shapley value, and the nucleolus. For instance, freire2014hybrid; nguyen2018sharing developed Benders decomposition and constraint-generation algorithms to solve mixed-integer linear programming formulations of the nucleolus problem, which aims to identify allocation rules within the core. Based on the selected allocation rule, the PM determines the market offer and subsequently distributes profits (or, equivalently, costs) to the WPPs. The operational framework implemented by the PM is illustrated in Figure 1.
@@
Although existing studies have investigated various allocation rules and computational techniques, they typically assume that the PM and WPPs have access to identical information and that WPPs unanimously agree on the submitted offers. In practice, however, diverse perspectives on future energy generation can lead to differing opinions between the PM and the WPPs regarding the appropriate offers. Specifically, each WPP typically uses a forecast model to estimate its own independent offer, as expressed in (6). Similarly, the PM employs a forecasting model $f$ with parameters $\phi$ to predict the aggregated offer based on the feature vector $\mathbf{x}_{t}$. This model is expressed as
$y^{\mathrm{F}}_{t+k,\mathrm{sum}}=\hat{y}_{t+k,\mathrm{sum}|t}=f(\mathbf{x}_{t};\phi).$ (11)
As the forecasts $\left\{\hat{y}_{t+k,\mathrm{sum}|t},\hat{y}_{t+k,1|t},\ldots,\hat{y}_{t+k,m|t}\right\}$ are generated independently by the PM and WPPs, they are likely to be incoherent. Thus, WPPs may disagree with the offer proposed by the PM, and furthermore, the forecasts cannot be directly used to determine each WPP’s share in the submitted offer. We address this challenge by developing a value-oriented forecast reconciliation approach in Section 3.
@@ 3.1 Problem Formulation
For ease of reference, the following concise notations are introduced: $\left[y_{t+k,\mathrm{sum}},y_{t+k,1},\ldots,y_{t+k,m}\right]^{\top}$ is represented by $\mathbf{y}_{t+k}$, while $\left[\hat{y}_{t+k,\mathrm{sum}|t},\hat{y}_{t+k,1|t},\cdots,\hat{y}_{t+k,m|t}\right]^{\top}$ is expressed as $\hat{\mathbf{y}}_{t+k|t}$, and referred to as the base forecast. As shown in Figure 1, the interaction between WPPs and the PM inherently establishes a two-level hierarchical structure. In the context of forecast reconciliation (athanasopoulos2023forecast), forecasts at the leaf nodes of the hierarchy are referred to as bottom-level forecasts, while forecasts at other nodes are called aggregated forecasts. In our specific WPP-PM hierarchy, we can decompose the vector $\mathbf{y}_{t}$ into $m$ bottom-level values $\mathbf{b}_{t}\in\mathbb{R}^{m}$ and an aggregated value $y_{t,\mathrm{sum}}$ such that $\mathbf{y}_{t}=\left[y_{t,\mathrm{sum}},\mathbf{b}_{t}^{\top}\right]^{\top}$. By defining a structural matrix $\mathbf{S}\in\{0,1\}^{(m+1)\times m}$, we obtain
$\mathbf{y}_{t}=\mathbf{S}\mathbf{b}_{t}\Leftrightarrow\begin{bmatrix}y_{t,\mathrm{sum}}\\ \mathbf{b}_{t}\end{bmatrix}=\begin{bmatrix}\mathbf{s}_{\mathrm{sum}}\\ \mathbf{I}_{m}\end{bmatrix}\mathbf{b}_{t},\ \forall t,$ (12)
where $\mathbf{s}_{\mathrm{sum}}=[1,1,\cdots,1]$ is an aggregation vector and $\mathbf{I}_{m}$ is the $m\times m$ identity matrix. Equivalently, (12) can be rewritten as
@@
where $\mathbf{B}=[1,-\mathbf{s}_{\mathrm{sum}}]$. Notably, the hierarchy does not have to be confined to just two levels, as it can accommodate nm>1n-m>1 aggregated series, meaning that $\mathbf{y}_{t}\in\mathbb{R}^{n}$. Correspondingly, the aggregation matrix is represented by $\mathbf{S}_{\mathrm{sum}}\in\{0,1\}^{(n-m)\times m}$, and $\mathbf{B}$ defined as $\mathbf{I}_{n-m}|-\mathbf{S}_{\mathrm{sum}}]$. With the hierarchical structure, the coherent forecast is formally defined as follows.
@@
A forecast $\hat{\mathbf{y}}_{t+k|t}$ is said to be coherent if $\mathbf{B}\hat{\mathbf{y}}_{t+k|t}=\mathbf{0}$.
As previously stated, the base forecast, $\hat{\mathbf{y}}_{t+k|t}$ is independently generated by WPPs and PM, and is likely not coherent. Therefore, it is essential to reconcile the forecasts before using them in market trading. Let $\tilde{\mathbf{y}}_{t+k|t}=\left[\tilde{y}_{t+k,\mathrm{sum}|t},\tilde{y}_{t+k,1|t},\ldots,\tilde{y}_{t+k,m|t}\right]^{\top}$ denote the reconciled forecasts, where
$\tilde{y}_{t+k,\mathrm{sum}|t}=\sum_{i=1}^{m}\tilde{y}_{t+k,i|t}.$ (14)
@@
$\tilde{\mathbf{y}}_{t+k|t}=f_{\mathrm{RE}}(\hat{\mathbf{y}}_{t+k|t}),$ (15)
where $f_{\mathrm{RE}}$ represents the reconciliation function. In other words, we transform incoherent base forecast $\hat{\mathbf{y}}_{t+k|t}$ into coherent forecast $\tilde{\mathbf{y}}_{t+k|t}$ via the reconciliation function. This process can be viewed as a special approach to forecast combination, integrating both direct predictions for each series and indirect forecasts derived from the structure (hollyman2021understanding).
After reconciliation, the PM then places offers in the market using the reconciled aggregate forecast, $\tilde{y}_{t+k,\mathrm{sum}|t}$. In parallel, the reconciled forecast for each WPP, $\tilde{y}_{t+k,i|t}$, is not submitted to the market; it serves instead to determine cost (or profit) allocation and is thus termed a pseudo-offer. In contrast to the independent offering outlined in Section 2, where all WPPs utilize an identical cost function, namely $c_{t+k}(\hat{y}_{t+k,i},y_{t+k,i})$, in this context, the WPPs encounter distinct cost functions. This variation arises because the WPPs are allocated as portions of the total balancing cost. Concretely, the allocated cost for WPP $i$ is calculated based on the reconciled forecasts $\tilde{\mathbf{y}}_{t+k|t}$ and realizations $\mathbf{y}_{t+k}$, which is represented as $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$. The detailed definition of $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ is postponed to Section 3.2. Correspondingly, the profit for WPP $i$ in the context of aggregated wind energy trading, is represented as $\rho_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$, which can be expressed as
$\rho_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})=\pi^{\mathrm{F}}_{t+k}y_{t+k,i}-c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k}).$ (16)
It is observed that, after reconciliation, the balancing cost of the aggregated forecast is always lower than or equal to the sum of costs of independent pseudo-offers, as described below.
Proposition 3.1.
Let $\left\{\tilde{y}_{t+k,1|t},\ldots,\tilde{y}_{t+k,m|t}\right\}$ be a set of $m$ independent pseudo-offers for the trading hour $t+k$. Then, it holds that,
$c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}})\leq\sum_{i=1}^{m}c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i}),$
where $\tilde{y}_{t+k,\mathrm{sum}|t}=\sum_{i=1}^{m}\tilde{y}_{t+k,i|t}$.
@@
Since the pseudo-offers $\tilde{y}_{t+k,i|t}$ are not actually submitted to the markets, the above proposition illustrates the potential cost implications of using reconciled forecasts as offers in the markets for any realization of the actual observations. Specifically, in the case of bottom-up reconciliation, each reconciled forecast $\tilde{y}_{t+k,i|t}$ coincides with the corresponding base forecast $\hat{y}_{t+k,i|t}$. This implies that the balancing cost of an aggregated offering can be lower than or equal to the sum of the costs incurred under independent offerings. Based on this, we state the following corollary.
Corollary 3.1.
After bottom-up reconciliation, the balancing cost of an aggregated offering is less than or equal to the sum of the costs incurred under independent offerings, for any realization of the actual observations.
Notably, coherence is often effectively ensured using existing forecast reconciliation methods. However, as noted by athanasopoulos2023evaluation, heterogeneous objectives sometimes arise across different levels of the hierarchy, with specific decisions tied to individual nodes. Despite this, most existing quality-oriented reconciliation approaches adopt a single loss function—typically the mean squared error—within the reconciliation process (athanasopoulos2023forecast). In contexts where agents possess heterogeneous cost functions, reconciliation naturally becomes a multi-objective optimization problem, highlighting the inherent challenge of reaching consensus on a suitable reconciliation function. Before presenting our proposed method, we outline several key assumptions, deferring detailed methodology to Section 3.2.
@@
In this context, we exclude sub-coalitions, as the forecasts for potential sub-coalitions are not available. Each WPP will decide whether to enter into an agency contract with the PM. Even if some WPPs choose to opt out of the contract, the PM can still manage the portfolio consisting of the remaining WPPs.
@@
This assumption is valid because WPPs are primarily focused on the individual benefits they derive from the coalition, rather than on the collective benefits of the coalition. As both $\alpha_{t+k}$ and $Y_{t,i}$ are time-varying, we introduce the following assumptions in line with besbes2015non.
Assumption 3.3.
We define the total variation of the sequence $\{\alpha_{t+k}\}$ as $V_{\alpha}$, given by
$V_{\alpha}=\sum_{t=1}^{T}|\alpha_{t+k}-\alpha_{t+k-1}|.$ (17)
We assume that $V_{\alpha}=o(T)$, implying that $\alpha_{t+k}$ varies slowly over time.
This assumption is practically reasonable, as electricity prices typically do not exhibit frequent abrupt changes.
Assumption 3.4.
For each distribution sequence $\{Y_{t,i}\}$ and $\{Y_{t,\mathrm{sum}}\}$, we define the total variation distance between consecutive distributions as $V_{Y,i}$, given by
$V_{Y,i}=\sum_{t=1}^{T}\frac{1}{2}\int|p_{t+k,i}(y)-p_{t+k-1,i}(y)|\mathrm{d}y.$ (18)
We assume that $V_{Y,i}=o(T)$, implying that $Y_{t,i}$ evolves slowly over time.
This assumption permits the time series to be non-stationary, provided that the underlying process exhibits only slow drift over time. Given that wind power time series are often quasi-stationary in practice, this assumption is also well justified.